OpenStudy (anonymous):

A bank says you can double your money in 10 years if you put $1,000 in a simple interest account. What annual interest rate does the bank pay?

OpenStudy (anonymous):

Not a friggin lot

OpenStudy (anonymous):

HOW MUCH IS THE SIMPLE INTEREST?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Sorry for the caps , But does it say anything about the percentage ?

OpenStudy (anonymous):

wut

OpenStudy (anonymous):

the % for the simple interest?

OpenStudy (anonymous):

do u know it

OpenStudy (anonymous):

I'm saying does the question say anything with a percent

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok tell me

OpenStudy (anonymous):

thats wwhat i want to know

OpenStudy (anonymous):

I was trying to find the formula but I really cant find it try clicking on members online and @ their username and they should be sure to help you with this

OpenStudy (mathmale):

A bank says you can double your money in 10 years if you put $1,000 in a simple interest account. What annual interest rate does the bank pay? Sounds like annual compounding to me (not simple interest). If the Principal is P, the interest rate r, the time (in years) t, then the Amount, A, is A=p(1+r)^t. So, since this fortunate investor ends up with twice as much money after 10 years, A=2P=P(1+r)^10.\[A=2P=P(1+r)^t\]

OpenStudy (whpalmer4):

No, it's a simple interest problem, with an interesting solution. \[I = Prt\]To double your money, \(I = P\), giving us\[P = Prt\]\[t = 10\]\[r =\]

OpenStudy (mathmale):

If \[2P=P(1+r)^{10},then. 2=(1+r)^{10}\]

OpenStudy (whpalmer4):

It graphically illustrates the power of compound interest vs. simple interest: the doubling time for compound interest at 7% is about 10 years. Look at how much more interest (higher rate) has to be paid to double your money in 10 years with simple interest!

OpenStudy (mathmale):

and the problem boils down to solving this equation for r. How? Take the logarithm (log or ln) of both sides. Try this. Ask if you need further guidance in solving for r.

OpenStudy (anonymous):

here are the options A1% B 1.25% C 10% D 12.5%

OpenStudy (whpalmer4):

Yeah, those answers are proof that the problem is looking for simple interest, as the 7% or so that would be the answer for compound interest doesn't appear, whereas the answer for simple interest does.

OpenStudy (mathmale):

(Face red) I overlooked the phrase, "simple interest," in the problem statement. Therefore, kudos to @whpalmer!

OpenStudy (whpalmer4):

I already answered this question for someone else last night, so I've had practice reading it :-)

OpenStudy (mathmale):

:-)

OpenStudy (anonymous):

wuts the answer

OpenStudy (mathmale):

We'd be happy to guide you towards finding the answer yourself. What do you need to know to be able to do that?

OpenStudy (anonymous):

dont know how

OpenStudy (mathmale):

You start with $1000. You double your money. How much interest did you earn?

OpenStudy (anonymous):

10% ??

OpenStudy (mathmale):

Hint: Your money doubles due to the addition of the interest earned. How much interest was earned, if the Amount in your bank account started at $1000, but is now $2000?

OpenStudy (anonymous):

12.5% ??

OpenStudy (whpalmer4):

@kjjk @mathmale is asking you for the amount (in dollars) of interest earned, not the rate...

OpenStudy (anonymous):

the question is asking for the rate

OpenStudy (whpalmer4):

I understand that, but we're trying to walk you through the entire process. You did ask for help, why not allow us to share our experience and knowledge?

OpenStudy (mathmale):

True, but there are several intermediate steps along the way towards finding the rate.

OpenStudy (anonymous):

im comfused

OpenStudy (mathmale):

You started with $1000; you ended up with $2000 after 10 years. How much interest did your money earn?

OpenStudy (anonymous):

1 sec

OpenStudy (mathmale):

Note: in this question I meant "an amount of money," not "interest rate." We'll calculate the interest rate shortly. But not yet.

OpenStudy (anonymous):

intrest rate is 50?

OpenStudy (mathmale):

Hint: Amount = Principal + Interest earned Here: $2000 = $1000 + ( ?? )

OpenStudy (anonymous):

1000

OpenStudy (mathmale):

Great. Thanks for your persistence. Then, as whpalmer4 has pointed out, i-Prt. Simple interest is equal to the product of the Principal, the interest rate and the time (in years). Since you have correctly identified the interest earned as being $1000, $1000 = P*r*t = $1000*r*10. Can you solve for r?

OpenStudy (mathmale):

\[1000=1000*r*10\]

OpenStudy (anonymous):

is the rate 12.5

OpenStudy (mathmale):

Please, would you show me the work you've done towards calculating the interest rate?

OpenStudy (mathmale):

\[1000=1000*r*10\] Please divide both sides of this equation by 1000 and write out your result.

OpenStudy (anonymous):

confused

OpenStudy (mathmale):

\[\frac{ 1000 }{1000 }=\frac{ 1000*r*10 }{ 1000 }\]

OpenStudy (mathmale):

this shows how I would divide both sides of the equation by 1000. Please simplify this result.

OpenStudy (anonymous):

so r=10

OpenStudy (whpalmer4):

No.

OpenStudy (mathmale):

1000/1000 on the left side results in 1 on the left side. 1000*r*10 on the right side, divided by 1000, results in 10*r on the right side. The equation now looks like 1 = 10*r

OpenStudy (anonymous):

1%=r

OpenStudy (mathmale):

Please solve this for r.

OpenStudy (mathmale):

If\[1=10r\]and we divide both sides by 10, we get\[\frac{ 1 }{ 10 }=r\]

OpenStudy (mathmale):

Please convert this to a decimal fraction.

OpenStudy (anonymous):

so it is 10%=r

OpenStudy (mathmale):

I'd prefer that you answer, "0.10," and only after that do you convert this decimal fraction to a percentage. 0.10*100% =?

OpenStudy (anonymous):

0.10*100% = 10

OpenStudy (mathmale):

Wouldn't that be 10%?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks makes senece now

OpenStudy (mathmale):

Congrats, @kjjk, you have solved this problem. The simple interest rate is 10%. It's very, very important that you know how to work with that equation for simple interest, i=Prt. This is why I would not answer "yes" or "no" when y ou asked whether the interest rate was 12.5%.

OpenStudy (anonymous):

thanks have to go

OpenStudy (mathmale):

Good luck! Hope to work with you again.

OpenStudy (anonymous):

bye